OpenStudy (anonymous):

Is 'Passing through' and "containing" the same thing when finding the equation of the line with the given properties?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

OK... Here is the problem to be solved: "Find the equation of the line with the given properties (-2,6) and (-1,8) Here is what I have done so far. My question to you so far is this: #1 am I correct up to the point I am at and from there, where do I go. ~ Below is my work Y - Y1 = M (X-X1) M=Y2-Y1 / X2-X1 (THESE ARE MY FORMULAS) M=8-6 / -1-(-2) = 2/1 Y-8 = 2/1 (X-1)

OpenStudy (anonymous):

you have done absolutely correct except for the last line Y-8 = 2/1 (X-(-1)) as th point is (-1,8)

OpenStudy (anonymous):

So so how you types it is how it should look?

OpenStudy (anonymous):

so the final ans would be (y-8)=2(x+1) or y-8=2x+2 or 2x-y +10=0

OpenStudy (anonymous):

you totally lost me. I was still trying to figure out the next step after Y-8=2/1 (x-(-1))

OpenStudy (anonymous):

can you back up a bit and walk me through the next few steps please?

OpenStudy (anonymous):

yup u ar right as (x-(-1)) = (x+1)

OpenStudy (anonymous):

and that's where I get lost. Y-8=2/1x+1 but I keep going blank as to how I get to the next step.

OpenStudy (anonymous):

Do I add 8 to both sides

OpenStudy (anonymous):

before this step ur correct Y-8 = 2/1 (X-1) if u take (x-1) it means that the line is passing through 1,8 hence instead of (x-1) it should be (x- (-1))=(x+1)

OpenStudy (anonymous):

So it now looks like: Y-8 = 2/1x + 1 my next step would be to add 8 to both sides? Y = 2/1x + 9

OpenStudy (anonymous):

here first multiply 2(x+1) as (y-8) = 2(x+1) and then simplify ...

OpenStudy (anonymous):

y-8=2x+2 or y=2x+2+8 or y=2x+10 2x-y +10=0

OpenStudy (anonymous):

So now I have Y= 2x + 10

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

You are awesomeness wrapped in a bow!!!!!!!!!!! Thank You Thank You Thank You